How do I find the inverse of: f(x)=2log2(x-5)? Please note that the "2" after log is in subscript. Please show steps.
Answer by stanbon(75887)

How do I find the inverse of: f(x)=2log2(x-5)? Please note that the "2" after log is in subscript. Please show steps.
-------------------------------------------
y = 2*log(base2)(x-5)
---
Interchange x and y:
x = 2*log(base2)(y-5)
Solve for "y":
log(base2)(y-5) = x/2
y-5 = 2^(x/2)
y = 2^(x/2) + 5
That is the inverse.
======================
Cheers,
Stan H.